OpenStudy (anonymous):

If given f(x) = x + 4 and g(x) = -2x to the second power + 3x -1 what is (f -g)(x) please help

OpenStudy (anonymous):

(f-g)(x) = f(x) - g(x) = x+4 - (2x^2 +3x-1) = x+4 -2x^2 -3x +1 = -2x^2 -2x +5

OpenStudy (anonymous):

Hope that helps :)